# EVF-SAM: Early Vision-Language Fusion for Text-Prompted Segment Anything Model

### v3

@@ 3.1 Motivation: SAM with Vision-Language Models
Early-fused architecture. As illustrated in Fig. 3 (b), we define the fusion for separately encoded single-modal prompts as ‘late fusion’, e.g., LLaVA [35]. Inspired by ViLT [42] and BEIT-3 [30], we further explore the ‘early-fusion’ paradigm for image and text modalities, which incorporates the cross-modal fusions within the encoder, e.g., [42, 30] perform cross-model fusions in the attention blocks. As shown in Fig. 3 (c), we leverage the ‘early-fusion’ vision-language model as the Multimodal Encoder to generate prompt embeddings for SAM. Tab. I shows that our investigation indicates that early-fusion outperforms late-fusion, i.e., 83.7 for BEIT-3 and 67.9 for CLIP. We believe the early-fused architecture is beneficial for encoding text prompts since the cross-modal fusions will further enhance the semantic representation for text embeddings. In addition, the text-to-image fusions guide the image branch to aggregate features which are aligned with text prompts, making the output embeddings more accurate for prompt SAM.
Encoder-based feature extractor. Recently, LISA [20, 21] and several LLM-based methods [27, 22] acquire the prompt embeddings for SAM with a special token through the auto-regressive generation. However, the uncontrolled length of the answering query introduces instability during both training and inference. Forcing the model to conform to a specific answering template can lead to language drift. In contrast, encoder-based architectures can maintain a consistent sequence length of inputs and outputs. Utilizing the encoder-based method not only offers convenience but also yields superior performance, i.e., 79.1 for LLaVA and 83.7 for BEIT-3. Notably, the encoder-based text-prompted SAM will reduce a massive computation burden compared to the LLM-based methods.
3.2 Architecture
@@
Multimodal Encoder. The Early Vision-Language Fused encoder adopts the input image and text and outputs fused multimodal embeddings. In EVF-SAM, we mainly adopt BEIT-3 [30] as the Multimodal Encoder, which formulates a multi-way transformer. The text is tokenized by XLMRobertaTokenizer [43] while the image is resized to $224^{2}$ and patched by a 1/16 convolution layer. Within each block of the encoder, the image and text tokens will be fused in the attention block and then fed into separate Feed-Forward Networks (FFN). We follow ViT [38] to retrieve the [CLS] token as the output multimodal embeddings.
Projector. Different foundation models tend to have different embedding dimensions (1024 for BEIT-3-Large, 768 for BEIT-3-Base, and 256 for SAM mask decoder). We adopt a simple MLP projector containing 2 Linear layers, activated by ReLU. In EVF-SAM, we do not design elaborate modules for better performance due to the following reasons: (1) the simple MLP is effective enough [35, 42], (2) using MLP is efficient for training and inference, and (3) the simple projector will have few impacts on the pre-trained knowledge of foundation models.
Adapted prompt encoder for SAM. SAM contains 3 main modules: (a) Image Encoder: a Vision Transformer [44] (ViT), extracting fine-grained feature maps from the input image. (b) Prompt Encoder: receiving interactive prompts and encoding them into hidden embeddings. (c) Mask Decoder: a lightweight mask generator to output the final masks based on previous embeddings. In EVF-SAM, we maintain the architecture of the image encoder and mask decoder while extending the prompt encoder to further gather the embeddings from the Multimodal Encoder. Specifically, the original prompt encoder encodes point or box prompts to sparse embeddings of $R^{B\times N\times D}$, where $B$, $N$, and $D$ refer to the batch size, number of points/boxes, and the embedding dimension, respectively. In EVF-SAM, the projected multimodal embeddings of $R^{B\times 1\times D}$ from the Multimodal Encoder will be concatenated to a zero-initialized sparse embeddings and then fed into the mask decoder.
@@ 3.3 Training
Template free. In most LLM-based frameworks, e.g., LISA [20, 21], Question-Answering templates are needed and designed to prompt the Large Multimodal Models (LMM) for the segmentation task, such as the instruction template “Can you segment {object} in the picture” and the answer template “It is [SEG].”. The absence of such templates can limit the effectiveness of pre-trained knowledge. Furthermore, the reliance on manually designed templates introduces uncertainty in training and can negatively impact inference results due to inconsistent question syntax from users. In contrast, EVF-SAM does not require pre-training on Question-Answering datasets, thus eliminating the need for templates. We simply adopt the expression phrases or sentences as input. This template-free approach simplifies training and inference.
Trainable modules. The Multimodal Encoder (EVF) is fully trainable during our training process, allowing it to learn how to generate multimodal embeddings tailored for SAM, which requires sufficient localization information for segmentation. For SAM, we keep the image encoder frozen during training while we enable training for the prompt encoder and mask decoder. Our experiments revealed that freezing the prompt encoder and mask decoder only leads to a minimal performance drop while maintaining SAM’s ability. We present details in Sec. III.
@@ 4.1 Datasets and Metrics
Datasets. We mainly conduct the experiments on RefCLEF [45], RefCOCO, RefCOCO+ [31, 45], and RefCOCOg [46, 47]. Specifically, RefCOCOg contains longer expressions which are manually annotated. Except for RefCOCO+, all datasets include geometric expression (e.g., ‘on the left’). Among different splits of testing datasets, ‘testA’ is human-centric, while ‘testB’ aims for common objects.
@@ 4.2 Implementation Details
Unless specified, we initialize the proposed EVF-SAM with the public weights of SAM-ViT-Huge [^1] SAM: https://github.com/facebookresearch/segment-anything [1] and BEIT-3-Large [^2] BEIT-3: https://github.com/microsoft/unilm [30]. All models are trained on 4 NVIDIA L40s GPUs with mixed precision. We adopt DeepSpeed [48] with ZeRO-2 for model parallel to optimize memory consumption. During training, the batch size of each GPU is 16 and we use gradient accumulation for 2 steps, therefore the total batch size per iteration is 128. We adopt AdamW [49] optimizer and set the initial learning rate to 1e-4 with a linear-decay schedule. We train all models for 15k iterations (nearly 1 day) and use the binary cross-entropy loss (BCE) and dice loss (the weight of both losses is 1.0).
TABLE II: Comparison of cIoU on different benchmarks between our proposed EVF-SAM and previous state-of-the-art methods. Bold: the best results. Underline: the second-best results. AVG represents the average metric across the eight RefCOCO-series benchmarks. We abbreviate the datasets: COCO (C), RefCOCO (RC), Object365 (O365), Video segmentation datasets (V), ADE20K (A), COCO-Stuff (CS), PACO-LVIS (PL), PASCAL-Part (PP), GranD (G), VOC2010 (VOC), MUSE (M), gRefCOCO (gRC), COCO-Interactive (CI), FSS-1000 (F), SA-1B (SA).
Method Foundation Models w/ SAM? Training Data RefCOCO () RefCOCO+ () RefCOCOg () AVG
@@
LAVT [15] BERT-B [36] (104M) RC, gRC 72.7 75.8 68.8 62.1 68.4 55.1 - - -
PolyFormer-L [16] BERT-B [36] (104M) RC, gRC 76.9 78.5 74.8 72.2 75.7 66.7 71.2 71.2 73.4
UNINEXT-H [19] BERT-B [36] (104M) O365, C, RC, V 82.2 83.4 81.3 72.5 76.4 66.2 74.4 76.4 76.6
UniLSeg-100 [18] CLIP-B [29] (63M) SA, RC, gRC 81.7 83.2 79.9 73.2 78.3 68.2 - - -
UniRef++-L [17] BERT-B [36] (104M) RC, F, V 79.1 82.1 77.5 68.4 74.0 61.5 71.4 72.8 73.4
LISA [20] Vicuna-7B [50] A, CS, RC, PL, PP 74.1 76.5 71.1 62.4 67.4 56.5 66.4 68.5 67.9
PixelLM [22] LLaMA2-13B [51] A, CS, RC, PL, M 73.0 76.5 68.2 66.3 71.7 58.3 69.3 70.5 69.2
GLaMM [27] Vicuna-7B [50] G, RC 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9 75.6
u-LLaVA [24] Vicuna-7B [50] A, CS, RC, PL, VOC 80.4 82.7 77.8 72.2 76.6 66.8 74.8 75.6 75.9
PSALM [25] Phi-1.5 [52] (1.3B) C, RC, CI 83.6 84.7 81.6 72.9 75.5 70.1 73.8 74.4 77.1
EVF-SAM BEIT-3 [30] (0.7B) RC 82.1 83.7 80.0 75.2 78.3 70.1 76.8 77.4 78.0
EVF-SAM-2 BEIT-3 [30] (0.7B) RC 82.7 84.1 80.0 76.3 80.1 71.8 77.0 78.4 78.8
@@ 4.3 Main Results
We mainly report the cIoU metric of RefCOCO-series benchmarks and compare our proposed EVF-SAM with recent state-of-the-art methods in Tab. II The upper part of Tab. II presents traditional methods based on text encoders. Despite their advantages in terms of fewer parameters and faster inference speeds, these methods either achieve less competitive results or require vast amounts of data due to their lack of integration with foundation models. The methods listed in the lower portion of Tab. II are based on Large Multimodal Models (LMMs), achieving state-of-the-art (SOTA) performance but require significant computational resources. Our EVF-SAM achieves the highest average cIoU score across all RES benchmarks, using only limited data and manageable computation costs. Specifically, our EVF-SAM achieves SOTA performance on RefCOCOg [46, 47], predicating a stronger capability for handling longer text prompts than previous LMM-based models, which is counter-intuitive while showing the great potential of vision-language models for understanding instructions. In addition, the early fusion between the input image and text prompts can generate more informative embeddings than independent encoders as discussed in Sec. 4. Furthermore, we incorporate the recently proposed SAM-2 [53] into EVF-SAM, and Tab. II shows that EVF-SAM-2 obtains significantly better performance on RefCOCO/+/g [31, 45], demonstrating the effectiveness of the proposed formulation of EVF-SAM.
@@ 4.4 Ablation Study
Multimodal Encoder and fusion methods. In Tab. III, we explore the effects of different Multimodal Encoders, e.g., CLIP, ViLT [42], and BEIT-3, and fusion methods, e.g., late fusion or early fusion. As shown in Tab. III, using a text-only encoder in EVF-SAM obtains limited segmentation performance on RefCOCO. Using Multimodal Encoders with both image and text inputs remarkably improves 4.5 cIoU, 4.6 cIoU, 4.0 cIoU, 1.0 cIoU, and 4.5 cIoU for CLIP-Large (OpenAI [^3] OpenAI: https://github.com/openai/CLIP), CLIP-Large (OpenCLIP [^4] OpenCLIP: https://github.com/mlfoundations/open_clip), CLIP-Huge (OpenCLIP), ViLT, and BEIT-3, respectively. It demonstrates the superiority of using multimodal prompts (text and input image) and showcases that the image embeddings will also provide useful guidance for SAM to segment objects accurately. We further evaluate the effects of early fusion on ViLT and BEIT-3, which adopts modality fusions in all self-attention layers. Specifically, we adopt two settings for BEIT-3 to analyze, e.g., fusions among former 12 layers ($L_{1}\sim L_{12}$), and fusions among all layers ($L_{1}\sim L_{24}$). Tab. III indicates that BEIT-3 with early fusion (fusing former 12 layers or fusing all 24 layers) significantly improves compared to late fusion or using text only. In addition, ViLT with early fusion also achieves 11.1 cIoU improvements compared to the baseline with text-only prompts, showing the effectiveness of early fusion and multimodal inputs for prompting SAM. Therefore, Tab. III demonstrates that (1) Multimodal Encoder with the input image and text and (2) early fusions between the image and text encoder are much effective for text-prompted SAM.
TABLE III: Ablation on fusion methods. We evaluate the performance of using different pre-trained Multimodal Encoders in EVF-SAM, e.g., CLIP from OpenAI [29] or OpenCLIP [54]. $L_{i}$ denotes the $i$-th layer in the BEIT-3 model (totally 24 layers for BEIT-3-Large). Half of the layers are activated to assess the impact of the modality fusion stage on model performance. $\dagger$: pre-trained models provided by OpenAI. $\ddagger$: pre-trained models provided by OpenCLIP.
Encoder Params Text Image Modality Fusion cIoU
@@
CLIP-Large 123M - 63.4
CLIP-Large 428M Late (Concat) 67.9
CLIP-Large 123M - 63.2
CLIP-Large 428M Late (Concat) 67.8
CLIP-Huge 302M - 64.2
CLIP-Huge 986M Late (Concat) 68.2
@@
ViLT 133M - 63.0
ViLT 136M Late (Concat) 64.0
ViLT 136M Early 75.3
BEIT-3-Large 370M - 65.1
BEIT-3-Large 673M Late (Concat) 69.6
BEIT-3-Large 673M Early ($L_{1}\sim L_{12}$) 82.2
BEIT-3-Large 673M Early ($L_{1}\sim L_{24}$) 83.7
Ablations on trainable modules. In Tab. IV, we evaluated the effects of fine-tuning (✓) or freezing ($\ast$) modules in the proposed EVF-SAM, i.e., the Multimodal Encoder, the prompt encoder, and the mask decoder. The image encoder of SAM is kept frozen during training. As Tab. IV shows, fine-tuning the Multimodal Encoder is crucial and it adapts the Multimodal Encoder to encode text and image inputs to multimodal representation for referring image segmentation. Notably, EVF-SAM can achieve competitive results with all modules of SAM kept frozen, and it can be seamlessly regarded as a strong extension for the original SAM, which simultaneously supports text prompts, box prompts and point prompts. Tab. IV Further fine-tuning the prompt encoder and mask decoder of SAM brings significant improvements.
TABLE IV: Ablations on trainable modules. We mainly evaluate the effects of fine-tuning or freezing the Multimodal Encoder, the prompt encoder and mask decoder of SAM. ‘✓’ denotes trainable, while $\ast$’ denotes frozen.
Multimodal Encoder Prompt Encoder Mask Decoder cIoU
@@
TABLE V: Ablations on multimodal feature representation. BEIT-3 contains two [CLS] tokens for visual and textual modalities. We also explore the effects of using AvgPool and combination between two modalities, same as late fusion.
Text-[CLS] Image-[CLS] Image-AvgPool Fusion cIoU
@@
Multimodal feature representation. In Tab. V, we explore the effects of using different multimodal features representations as prompts for SAM. Specifically, we adopt different outputs of the Multimodal Encoder: (a) the image [CLS] token, (b) the AvgPool over image tokens, and (c) the text [CLS] token. Tab. V shows that using image [CLS] token is more effective while combining image and text tokens through concatenation leads to a performance drop.
Effects of extra semantic data. While focusing on referring expression segmentation tasks, we find that mainstream works (e.g., u-LLaVA [24], PSALM [25]) emphasize model ability on various extra tasks. We also prove that the effectiveness of our EVF-SAM extends to other tasks as well. We introduce some extra semantic segmentation datasets (ADE20k [55], Mapillary [56]) to proceed with joint training. We do not include COCO-Stuff [57] to avoid data leakage with RefCOCO/+/g. We report comparison results in Tab. VI. We find our EVF-SAM able to exhibit exceptional performance across multiple tasks simultaneously. Specifically, (a) we find performance gain on the RefCOCO+ benchmark when introducing ADE20k. We hypothesize the reason lies in the analogous data composition between RefCOCO+ and ADE20k; (b) We observe performance degradation on RefCOCO and RefCOCOg. This is due to the differences in text expression syntax between semantic segmentation data and RES data; (c) Our experiments indicate that the zero-shot inference capability of EVF-SAM can be augmented by incorporating additional semantic segmentation data.
TABLE VI: Results of adding extra semantic data. means zero-shot results. The reported ADE20k results are evaluated on the validation set using the cIoU metric.
ADE20k Mapillary RefCOCO RefCOCO+ RefCOCOg ADE20k
val testA testB val testA testB val test val
82.1 83.7 80.0 75.2 78.3 70.1 76.8 77.4   54.2
81.7 83.6 80.3 75.4 78.4 71.3 75.5 77.6 75.9
81.9 83.5 80.3 75.1 78.0 70.8 75.3 77.4   59.6
81.8 83.4 79.7 75.6 78.0 70.7 75.8 76.9 76.1
Effects of Different Foundation Models. In Tab. VII, we explore the effects of using different foundation models in EVF-SAM. For the Multimodal Encoder, we adopt CLIP-Large (only text encoder), ViLT, BEIT-3-Large, and BEIT-3-Base. We also modify EVF-SAM with Efficient-SAM [2] to formulate a lighter version, which reduces 600M parameters compared to SAM-H. As shown in Tab. VII, EVF-SAM with BEIT-3-Base brings a severe performance drop which indicates a better Multimodal Encoder leads to better prompts for SAM. Remarkably, Tab. VII shows a negligible difference between Efficient-SAM-S and SAM-H in EVF-SAM, which demonstrates the effectiveness of Efficient-SAM and also indicates that EVF-SAM performs well for different SAM variants. In addition, it also provides insights about designing text-prompted SAMs for future research, e.g., developing a larger and better Multimodal Encoder is more important to empower SAM with text-prompted abilities.
TABLE VII: Comparison of effects of different foundation models. AVG represents the average metric across the eight RefCOCO-series benchmarks.

### v4

@@ 3.2 Architecture
Multimodal Encoder. The Early Vision-Language Fused encoder adopts the input image and text and outputs fused multimodal embeddings. In EVF-SAM, we mainly adopt BEIT-3 (Wang et al., 2022a) as the Multimodal Encoder, which formulates a multi-way transformer. The text is tokenized by XLMRobertaTokenizer (Conneau et al., 2019) while the image is resized to $224^{2}$ and patched by a 1/16 convolution layer. Within each block of the encoder, the image and text tokens will be fused in the attention block and then fed into separate Feed-Forward Networks (FFN). We follow ViT (Dosovitskiy et al., 2020) to retrieve the [CLS] token as the output multimodal embeddings.
Projector. Different foundation models tend to have different embedding dimensions (1024 for BEIT-3-Large, 768 for BEIT-3-Base, and 256 for SAM mask decoder). We adopt a simple MLP projector containing 2 Linear layers, activated by ReLU. In EVF-SAM, we do not design elaborate modules for better performance due to the following reasons: (1) the simple MLP is effective enough (Liu et al., 2023b; Kim et al., 2021), (2) using MLP is efficient for training and inference, and (3) the simple projector will have few impacts on the pre-trained knowledge of foundation models.
Adapted prompt encoder for SAM. SAM contains 3 main modules: (a) Image Encoder: a Vision Transformer (Li et al., 2022) (ViT), extracting fine-grained feature maps from the input image. (b) Prompt Encoder: receiving interactive prompts and encoding them into hidden embeddings. (c) Mask Decoder: a lightweight mask generator to output the final masks based on previous embeddings. In EVF-SAM, we maintain the architecture of the image encoder and mask decoder while extending the prompt encoder to further gather the embeddings from the Multimodal Encoder. Specifically, the original prompt encoder encodes point or box prompts to sparse embeddings of $R^{B\times N\times D}$, where $B$, $N$, and $D$ refer to the batch size, number of points/boxes, and the embedding dimension, respectively. In EVF-SAM, the projected multimodal embeddings of $R^{B\times 1\times D}$ from the Multimodal Encoder will be concatenated to a zero-initialized sparse embeddings and then fed into the mask decoder.
@@ 3.3 Training
Instruction template-free. In most LLM-based frameworks, e.g., LISA (Lai et al., 2023; Yang et al., 2023), instruction templates are required to prompt the Large Multimodal Models (LMM) for the segmentation task, e.g., ‘Can you segment {object} in the picture’ with answer ‘It is [SEG].’. Removing instruction templates will affect the performance of LMMs, therefore, users need to follow the corresponding templates for referring image segmentation. In contrast, EVF-SAM does not require pre-training on QA (question-answering) datasets, thus eliminating the need for instruction templates. We adopt the expression phrases or sentences as input. This template-free approach simplifies training and inference.
Trainable modules. The Multimodal Encoder (EVF) is fully trainable during our training process, allowing it to learn how to generate multimodal embeddings tailored for SAM, which requires sufficient localization information for segmentation. For SAM, we keep the image encoder frozen during training while we enable training for the prompt encoder and mask decoder. Our experiments revealed that freezing the prompt encoder and mask decoder only leads to a minimal performance drop while maintaining SAM’s ability. We present details in Sec. 3.
Unified training with multi-tasks. To further enhance the generic multi-task segmentation capabilities of EVF-SAM, including semantic segmentation and fine-grained part segmentation, we present a unified training strategy for EVF-SAM with diverse training datasets, such as ADE20K (Zhou et al., 2017b), PartImageNet (He et al., 2022) and PASCAL-Part (Chen et al., 2014).
However, we observe a performance degradation when simply mixing the training data of referring and semantic segmentation (shown in Tab. 7 of the Appendix), which can be attributed to the semantic conflict among different tasks, as discussed in UniLSeg (Liu et al., 2023e). To alleviate the aforementioned conflicts, we leverage a special text token [semantic] and input ‘[semantic]{category}’ for semantic/part segmentation.
@@ 4.1 Datasets and Metrics
Datasets. We mainly conduct the experiments on RefCLEF (Kazemzadeh et al., 2014), RefCOCO, RefCOCO+ (Yu et al., 2016; Kazemzadeh et al., 2014), and RefCOCOg (Nagaraja et al., 2016; Mao et al., 2016). Specifically, RefCOCOg contains longer expressions which are manually annotated. Except for RefCOCO+, all datasets include geometric expression (e.g., ‘on the left’). Among different splits of testing datasets, ‘testA’ is human-centric, while ‘testB’ aims for common objects.
Extra training datasets. To further enhance the versatility of EVF-SAM, we employ multi-task unified training by expanding the training datasets by introducing Objects365 (Shao et al., 2019), ADE20K (Zhou et al., 2017b), PASCAL-Part (Chen et al., 2014), PartImageNet (He et al., 2022), and HumanParsing (Liang et al., 2015b). Therefore, EVF-SAM can handle various granularity of text-prompted segmentation, e.g., semantic-level, instance-level, and part-level segmentation. We refer the readers to the appendix for more details about training with the extra multi-task datasets.
@@ 4.2 Implementation Details
Unless specified, we initialize the proposed EVF-SAM with the public weights of SAM-ViT-Huge  [^1] SAM: https://github.com/facebookresearch/segment-anything (Kirillov et al., 2023) and BEIT-3-Large  [^2] BEIT-3: https://github.com/microsoft/unilm (Wang et al., 2022a). All models are trained on 4 NVIDIA L40s GPUs with mixed precision. We adopt DeepSpeed (Song et al., 2023) with ZeRO-2 for model parallel to optimize memory consumption. During training, the batch size of each GPU is 16 and we use gradient accumulation for 2 steps, therefore the total batch size per iteration is 128. We adopt AdamW (Loshchilov & Hutter, 2017) optimizer and set the initial learning rate to 1e-4 with a linear-decay schedule. We train all models for 15k iterations (nearly 1 day) and use the binary cross-entropy loss (BCE) and dice loss (the weight of both losses is 1.0).
Table 2: Comparison of cIoU on different benchmarks between our proposed EVF-SAM and previous state-of-the-art methods. Bold: the best results. Underline: the second-best results. AVG represents the average metric across the eight RefCOCO-series benchmarks. We abbreviate the datasets: COCO (C) (Lin et al., 2014), RefCOCO (RC) (Yu et al., 2016; Nagaraja et al., 2016; Mao et al., 2016; Kazemzadeh et al., 2014), Objects365 (O) (Shao et al., 2019), Video segmentation datasets (V), ADE20K (A) (Zhou et al., 2017a; 2019), COCO-Stuff (CS) (Caesar et al., 2018), PACO-LVIS (PL) (Ramanathan et al., 2023), PASCAL-Part (PP) (Chen et al., 2014), GranD (G) (Rasheed et al., 2023), PASCAL VOC2010 (PV) (Everingham et al., 2010), MUSE (M) (Ren et al., 2023), gRefCOCO (gRC) (Liu et al., 2023a), COCO-Interactive (CI) (Zhang et al., 2024), FSS-1000 (F) (Li et al., 2020), SA-1B (SA) (Kirillov et al., 2023), PartImageNet (PIN) (He et al., 2022), HumanParsing (HP) (Liang et al., 2015b; a), GoldG (GG) (Kamath et al., 2021).
Method Text Prompt Encoder SAM? Training Data RefCOCO RefCOCO+ RefCOCOg AVG
@@
[…]
EVF-SAM BEIT-3 RC 82.1 83.7 80.0 75.2 78.3 70.1 76.8 77.4 78.0
EVF-SAM BEIT-3 RC, O, A, PP, PIN, HP 82.4 84.2 80.2 76.5 80.0 71.9 78.2 78.3 79.0
@@ 4.3 Main Results
We mainly report the cIoU metric of RefCOCO-series benchmarks and compare our proposed EVF-SAM with recent state-of-the-art methods in Tab. 2 The upper part of Tab. 2 presents traditional methods based on text encoders. Despite their advantages in terms of fewer parameters and faster inference speeds, these methods either achieve less competitive results or require vast amounts of data due to their lack of integration with foundation models. The methods listed in the lower portion of Tab. 2 are based on Large Multimodal Models (LMMs), achieving state-of-the-art (SOTA) performance but require significant computational resources. Our EVF-SAM achieves the highest average cIoU score across all RES benchmarks, using only limited data and manageable computation costs. Specifically, our EVF-SAM achieves SOTA performance on RefCOCOg (Nagaraja et al., 2016; Mao et al., 2016), predicating a stronger capability for handling longer text prompts than previous LMM-based models, which is counter-intuitive while showing the great potential of vision-language models for understanding instructions. In addition, the early fusion between the input image and text prompts can generate more informative embeddings than independent encoders as discussed in Sec. 1.
@@ 4.4 Ablation Study
Table 3: Ablation on fusion methods. We evaluate the performance of using different pre-trained Multimodal Encoders in EVF-SAM, e.g., CLIP from OpenAI (Radford et al., 2021) or OpenCLIP (Ilharco et al., 2021). $L_{i}$ denotes the $i$-th layer in the BEIT-3 model (totally 24 layers for BEIT-3-Large). Half of the layers are activated to assess the impact of the modality fusion stage on model performance. $\dagger$: pre-trained models provided by OpenAI. $\ddagger$: pre-trained models provided by OpenCLIP.
Encoder Params Text Image Modality Fusion RefCOCO RefCOCO+ RefCOCOg AVG
val testA testB val testA testB val test
@@
CLIP-Large 123M - 61.0 63.4 59.9 43.1 45.9 40.6 48.9 49.6 51.6
CLIP-Large 428M Late (Concat) 67.4 68.9 64.4 50.5 54.6 46.7 55.1 56.2 58.0
CLIP-Large 123M - 60.8 63.2 59.0 42.9 46.4 39.2 49.2 50.5 51.4
CLIP-Large 428M Late (Concat) 66.1 67.8 63.1 49.8 51.9 44.1 54.1 55.0 56.5
CLIP-Huge 302M - 61.7 64.2 60.1 44.2 47.8 40.2 49.6 50.9 52.3
CLIP-Huge 986M Late (Concat) 66.3 68.2 64.3 49.8 53.5 45.1 55.4 56.7 57.4
@@
ViLT 133M - 61.0 63.0 60.0 42.5 45.4 39.5 49.3 49.5 51.3
ViLT 136M Late (Concat) 61.4 64.0 59.6 42.8 46.4 40.1 49.5 50.0 51.7
ViLT 136M Early 73.9 75.3 70.9 61.1 64.4 55.2 65.1 66.8 66.6
BEIT-3-Large 370M - 61.6 65.1 59.4 44.0 47.6 40.6 49.5 50.8 52.3
BEIT-3-Large 673M Late (Concat) 67.7 70.2 65.4 51.1 55.0 46.9 57.2 57.0 58.8
BEIT-3-Large 673M Early ($L_{1}\sim L_{12}$) 80.6 82.2 78.8 72.4 75.7 66.7 73.7 75.0 75.6
BEIT-3-Large 673M Early ($L_{1}\sim L_{24}$) 82.1 83.7 80.0 75.2 78.3 70.1 76.8 77.4 78.0
Multimodal Encoder and fusion methods. In Tab. 3, we explore the effects of different Multimodal Encoders, e.g., CLIP, ViLT (Kim et al., 2021), and BEIT-3, and fusion methods, e.g., late fusion or early fusion. As shown in Tab. 3, using a text-only encoder in EVF-SAM obtains limited segmentation performance on RefCOCO. Using Multimodal Encoders with both image and text inputs remarkably improves 4.5 cIoU, 4.6 cIoU, 4.0 cIoU, 1.0 cIoU, and 4.5 cIoU for CLIP-Large (OpenAI  [^3] OpenAI: https://github.com/openai/CLIP), CLIP-Large (OpenCLIP  [^4] OpenCLIP: https://github.com/mlfoundations/open_clip), CLIP-Huge (OpenCLIP), ViLT, and BEIT-3, respectively. It demonstrates the superiority of using multimodal prompts (text and input image) and showcases that the image embeddings will also provide useful guidance for SAM to segment objects accurately. We further evaluate the effects of early fusion on ViLT and BEIT-3, which adopts modality fusions in all self-attention layers. Specifically, we adopt two settings for BEIT-3 to analyze, e.g., fusions among former 12 layers ($L_{1}\sim L_{12}$), and fusions among all layers ($L_{1}\sim L_{24}$). Tab. 3 indicates that BEIT-3 with early fusion (fusing former 12 layers or fusing all 24 layers) significantly improves compared to late fusion or using text only. In addition, ViLT with early fusion also achieves 11.1 cIoU improvements compared to the baseline with text-only prompts, showing the effectiveness of early fusion and multimodal inputs for prompting SAM. Therefore, Tab. 3 demonstrates that (1) Multimodal Encoder with the input image and text and (2) early fusions between the image and text encoder are much effective for text-prompted SAM.
Ablations on trainable modules. In Tab. 5, we evaluated the effects of fine-tuning (✓) or freezing ($\ast$) modules in the proposed EVF-SAM, i.e., the Multimodal Encoder, the prompt encoder, and the mask decoder. The image encoder of SAM is kept frozen during training. As Tab. 5 shows, fine-tuning the Multimodal Encoder is crucial and it adapts the Multimodal Encoder to encode text and image inputs to multimodal representation for referring image segmentation. Notably, EVF-SAM can achieve competitive results with all modules of SAM kept frozen, and it can be seamlessly regarded as a strong extension for the original SAM, which simultaneously supports text prompts, box prompts and point prompts. Tab. 5 Further fine-tuning the prompt encoder and mask decoder of SAM brings significant improvements.
Table 4: Ablations on trainable modules. We mainly evaluate the effects of fine-tuning or freezing the Multimodal Encoder, the prompt encoder and mask decoder of SAM. ‘✓’ denotes trainable, while $\ast$’ denotes frozen.
Multimodal Enc. Prompt Enc. Mask Dec. cIoU
@@
Table 5: Ablations on multimodal feature representation. BEIT-3 contains two [CLS] tokens for visual and textual modalities. We also explore the effects of AvgPool and late fusion between two modalities.
[…]
@@
Multimodal feature representation. In Tab. 5, we explore the effects of using different multimodal features representations as prompts for SAM. Specifically, we adopt different outputs of the Multimodal Encoder: (a) the image [CLS] token, (b) the AvgPool over image tokens, and (c) the text [CLS] token. Tab. 5 shows that using image [CLS] token is more effective while combining image and text tokens through concatenation leads to a performance drop.
Effects of Different Foundation Models. In Tab. 6, we explore the effects of using different foundation models in EVF-SAM. For the Multimodal Encoder, we adopt CLIP-Large (only text encoder), ViLT, BEIT-3-Large, and BEIT-3-Base. We also modify EVF-SAM with Efficient-SAM (Xiong et al., 2023) to formulate a lighter version, which reduces 600M parameters compared to SAM-H. As shown in Tab. 6, EVF-SAM with BEIT-3-Base brings a severe performance drop which indicates a better Multimodal Encoder leads to better prompts for SAM. Remarkably, Tab. 6 shows a negligible difference between Efficient-SAM-S and SAM-H in EVF-SAM, which demonstrates the effectiveness of Efficient-SAM and also indicates that EVF-SAM performs well for different SAM variants. In addition, it also provides insights about designing text-prompted SAMs for future research, e.g., developing a larger and better Multimodal Encoder is more important to empower SAM with text-prompted abilities.
Table 6: Comparison of effects of different foundation models. AVG represents the average metric across the 8 RefCOCO-series benchmarks.